# Frontier AI’s Impact on the Cybersecurity Landscape

### v2

@@ 4.1.1 Qualitative analysis
Defense step 4&5: Remediation development and deployment. Research demonstrates frontier AI’s capability to automatically generate security vulnerability patches [172, 177, 159, 3, 332, 197, 218, 263, 321, 55]. To enhance patching effectiveness, researchers have explored various prompt strategies and built AI agents where LLMs retrieve contextual information (e.g., caller functions of vulnerable code) [34, 177]. The AIxCC competition demonstrates real-world applications, including patching an Off-by-One bug in SQLite3 [119]. However, no specific work exists on AI-facilitated remediation deployment.
Future impact. We envision that specialized reasoning models can generate better patches for complex vulnerabilities. Early research showed that fine-tuning LLMs through RL or supervised learning can enhance bug-fixing capabilities for non-security bugs [309, 144]. We can also build more advanced agents integrated with traditional program analysis techniques to generate, validate, and refine patches. Recent research has demonstrated the effectiveness of this approach for logical and non-security bugs [171, 312, 304]. It is also critical to provide formal guarantees for LLM-generated patches, especially for security vulnerabilities [46].
Remediation deployment remains a lengthy process requiring substantial manual effort [42, 66], e.g., hospital systems currently average 471 days for remediation deployment [42]. As shown in Figure 4, new patches must undergo heavy testing to ensure reliability and correctness and need to resolve any potential conflicts with their dependencies. We envision that frontier AI can also accelerate and automate remediation deployment, given they can accelerate general software development (e.g., Copilot [103], Cursor [65]). For example, code generation models can generate testing cases for functionality and security testing [237, 54, 334]. LLMs can correct configuration errors (e.g., Hadoop) and set up software environments (e.g., Linux server) [250, 38]. As AI models continuously improve their ability to understand and analyze large code projects, they could be applied to identify potential conflicts in dependent packages and even fix such conflicts. These capabilities of frontier AI can largely accelerate the remediation deployment and reduce the time window during which systems remain vulnerable.
TABLE III: Overview of existing defense benchmarks. “Vul. D” and “Att. D” refer to vulnerability detection and attack detection. “RT” is root cause and “RE” is reverse engineering. “Re. dev.” and “Re. dep.” stands for remediation development and remediation deployment. The performances of the SOTA LLM models on these benchmarks are shown in Table VII.
                                      Defense steps Quantity Notable examples Key limitations
1 Pen. test Few AutoPenBench [101] Limited attacks, target systems, metrics,
Vul. D Many PrimeVul [80],FuzzBench [199], OSS-Fuzz [247] limited complexity and diversity, low-quality labels
2 Att. D Many IDS2018 [252], Drebin [19], BODMAS [315] Low-quality data and labels, lack OOD/hidden tests
3 PoC & RT Few CRUXEval [108] Low complexity, no benchmark for root cause,
Rev. Eng. Many ByteWeight [25], ReSym [305] data leakage, incomplete task coverage
4 Rem. Dev. Few SWE-bench [147] Noisy, no benchmark for security bugs, limited programming languages (PLs)
5 Rem. Dep. None None N/A
4.1.2 Quantitative analysis
Table III shows an overview of existing defense benchmarks, where most benchmarks focus on vulnerability detection and attack detection, while other defense steps lack high-quality benchmarks. Table VII details SOTA LLMs’ performance on these benchmarks.
Defense Step 1: Proactive testing. Benchmarking penetration testing is related to attack benchmarking. As discussed in Section 3.1, penetration testing benchmarks remain at an early stage, where AutoPenBench [101] is the most recent benchmark enabling end-to-end penetration testing with dynamic metrics.
There are many vulnerability detection benchmarks, including static benchmarks with labeled code snippets [217, 80, 277, 47, 275, 284, 118, 212] and dynamic benchmarks with fuzzing instrumentation [247, 199]. The latest static benchmark, PrimeVul [80], contains vulnerable C/C++ and Python code covering over 100 CWEs, showing that SOTA LLMs achieve only $\sim$30F1 score. PrimeVul shares limitations with other static benchmarks, most notably low-quality labels. These include inaccurate labels caused by missing context (snippets only vulnerable under specific conditions) and incomplete labels (snippets with multiple vulnerabilities but only one labeled). In addition, existing benchmarks mostly cover snippets with only one function selected from limited projects. For dynamic datasets, OSS-Fuzz [247] is the largest fuzzing benchmark, and it is actively maintained and updated. FuzzBench [199] is a smaller but less noisy benchmark widely used in research papers to test novel fuzzing methods. These benchmarks are also used to test AI-facilitated fuzzing approaches, which achieved improvements over traditional fuzzers [256].
Future efforts. For penetration testing, future works should expand attack step coverage, diversify target systems, and implement more comprehensive evaluation metrics. For vulnerability detection, we should enhance label quality through human validation and include complex code snippets from diverse projects.
Defense Step 2: Attack detection. Many benchmarks exist for network intrusion detection and malware detection [252, 19, 315]. Recent research [18] highlights these benchmarks’ limitations in data quality (duplicated data, shortcuts) and label accuracy. Foundation model-driven detection shows very high accuracy on these benchmarks [117, 178]. Besides attack detections, CTIBench [8] designed tasks for LLMs to pinpoint the specific CWEs and attack techniques for given attack instances
Future efforts. Future work can improve data and label quality while incorporating more complex attack instances. Since existing AI-based detection lacks generalizability, benchmarks can include OOD testing sets with unseen attack types and drift patterns. Moreover, dynamic benchmark updates are essential to incorporate emerging attacks.
Defense Step 3: Triage & forensic. CRUXEval [108] benchmarks PoC generation for vulnerable C/C++ codes with overflow bugs, showing that GPT-4o achieves 75% pass@1 success rate.
Future efforts. We can include more vulnerabilities and cover more programming languages for PoC generation benchmarks. While no benchmark is publicly available for comprehensively evaluating AI-driven reverse engineering, most AI-based methods released their own datasets [257, 114, 305]. These datasets could potentially be merged to create a unified benchmark, covering tasks from function boundary identification to variable or data structure recovery. There is no public benchmark for root causes, and constructing such benchmarks is challenging because ground truth is hard to determine. One possible way is to extract samples from reproducible vulnerability datasets (e.g., ARVO [193]) and craft ground truth root causes via crowdsourcing.
Defense Step 4&5: Remediation development and deployment. Given that AI is not yet widely involved in these steps, there are not many benchmarks. The most popular benchmark is SWE-bench [147], which extracts issues from multiple Github Python projects. To fix the issues, we need to build AI agents that can analyze and locate the issues, generate patched code, and validate the generated patches. SOTA multi-agent systems with OpenAI o3 and Claude 3.5 Sonnet can resolve more than 55% issues in the SWE-bench-verified benchmark. Recent works extended the SWE-bench to multi-modal settings [312] and the Java language [325].
Future efforts. Given that the SWE-bench focuses on non-security bugs, it is necessary to construct benchmarks with security bugs covering multiple languages, especially C/C++. The procedure can be similar to root cause where we start from datasets with reproducible vulnerabilities. The main effort will be creating execution environments, crafting ground truth patches, and defining proper evaluation methods (e.g., writing test cases for dynamic testing).
4.1.3 Facilitate human defenders and SOC
Frontier AI can also improve the productivity of human defenders. Human and AI collaborative defenses could draw a balance between defense efficacy, robustness, and cost efficiency. For example, during proactive testing and attack detections, we can use AI to filter out the most obvious benign cases, which significantly reduces human efforts. AI can also propose initial PoCs, patches, and configurations, such that a human only needs to examine and fix these candidates rather than starting from scratch. Frontier AI is also enhancing Security Operations Centers (SOCs) by improving threat detection, incident response, and operational efficiency [95, 83]. Recent progresses [116, 259] show that AI-driven tools can autonomously investigate and respond to security alerts, enabling SOCs to analyze increasing amounts of security data and identify more threats in real-time. The continuous advancements in AI-driven tools are expected to enhance the productivity and scalability of SOCs [157].
@@ 4.2 Defenses with Provable Guarantees
Secure-by-design and secure-by-construction defenses provide formal security guarantees. Formal verification proves whether systems satisfy specific properties, with tools developed for various systems [246, 109, 164, 125, 48, 31], including seL4 [246] and CertiKOS [109] for OS kernels, CompCert [164] for compilers, IronFleet [125] for distributed systems, and EasyCrypt for cryptographic proofs [27]. This process is labor-intensive; seL4 required tens of proof-engineer years [129, 130]. Failed verifications require expert refinement, and software updates demand rewriting proofs, significantly increasing human effort.
Recent research explores frontier AI to improve formal verification’s automation and scalability. Some works enhanced interactive theorem proving [138, 322, 10, 314], e.g., GamePad [138] uses deep learning to synthesize Coq proofs. Others focus on automated program verification, where the typical solution is to model programs in domain-specific formal languages (DSLs), formulate security properties as specifications, and use constraint solvers for verification. Research works used LLMs to generate program invariants, lifting a program to a DSL or supplementary text hints [219, 150, 32, 300, 185]. For example, Loopy [150] leveraged LLMs to generate candidate loop invariants and refined invariants via an SMT solver. Lemur [300] used LLMs to generate high-level program invariants as sub-goals, which are then checked by solvers. LLMLift [32] combined LLMs with verified lifting tools to provide verified code transpilation from the source to DSLs, where LLMs generate lifted codes and proofs for functional equivalence. DafnyBench [185] evaluates SOTA LLMs in generating hints for the Dafny verifier. Other works improve solvers, including using traditional deep learning to enhance SAT and SMT solvers [24, 173] and leveraging LLMs for solver testing and fuzzing [268]. Section 7.2 shows our recommendations for improving formal verification with frontier AI.
@@ 4.3 Defense For Humans
Frontier AI enhances defenses against human-targeted attacks. LLMs show promise in detecting social bots [94], and AI has been deployed for real-world fraud detection [166, 326, 78]. Research demonstrates multi-modal agents capable of detecting phishing websites [39]. AI-driven solutions are increasingly being applied to real-world scenarios. For example, one company deployed an AI-based honeypot to combat vishing, using a chatbot with a “grandma” voice to occupy fraudsters and waste their time [198]. AI-based deepfake detection continues as an active research area [229], with exploration of LLMs’ capabilities in identifying fake images [6]. Another significant direction focuses on AI-powered misinformation detection [14, 79, 180], often employing LLMs to identify LLM-generated content. Watermarking techniques represent another promising detection approach [333, 214, 179, 335, 110].
@@
Despite these active research efforts, Figure 3 shows current defensive techniques struggling to match sophisticated attacks. Research reveals significant limitations in AI systems’ defensive capabilities. For example, studies examining GPT-4’s ability to detect fake images [6] concluded the model lacks reliability and shows limitations with complex tasks. While many papers propose AI-powered defenses against deepfakes and social bots, most address traditional attack scenarios rather than novel or sophisticated attacks enabled by frontier AI. We need studies on adversarial dynamics between defensive and attack AI systems—for example, how defensive AI can detect deepfakes or social bots generated by attack AI. This gap underscores the urgent need for research developing AI-driven defenses against AI-driven human-targeted attacks.
Future impact. As frontier AI evolves, its defensive capabilities will improve significantly. Future systems can detect and prevent attacks in real-time by analyzing patterns across multiple modalities simultaneously. These advanced systems could predict and preempt sophisticated social engineering attempts, deepfakes, and fraud schemes before reaching their targets. In addition, an increasing number of AI systems will be deployed as decoys [198, 230], posing as vulnerable targets to deceive attackers and exhaust their resources. By integrating frontier AI, authentication mechanisms can become more robust and adaptive to emerging threats, including sophisticated watermarking techniques and detection methods designed to counter evolving attack strategies.
However, this defensive potential faces a critical challenge: the inherent arms race between offensive and defensive capabilities suggests that as defensive systems become more sophisticated, attackers will likely develop increasingly advanced countermeasures using the same frontier AI technologies. We discuss this arms race between attacks and defenses in the next section.
Key takeaways. Qualitative: Frontier AI has been utilized in proactive testing and attack detection but remains underexplored in vulnerability triage and remediation. Existing AI-driven detections suffer from high false positive rates, limited generalizability, and insufficient robustness. Early AI-driven triage and patch tools cannot beat heuristic-based methods and manual analysis, nor can they provide formal correctness and security guarantees. Quantitative: Benchmarks show frontier AI outperforms traditional static detection methods and modestly improves fuzzing for vulnerability and attack detection. Future benchmarks for proactive testing and attack detection can enhance data quality and label accuracy, and include hidden OOD test sets. More comprehensive benchmarks are needed for other steps. Prediction: By training with high-quality datasets with rich program artifacts and stronger safety alignments, AI-driven detections can have better reasoning capabilities and improvements in generalizability and robustness. AI agents, integrating AI-driven planning and generation with program analysis tools, will harness the strengths of learning-based and program analysis-based approaches, potentially leading to more effective and automated triage and remediation. Systems with provable guarantees: Early attempts leveraged frontier AI for theorem proving and program verification. Frontier AI advances formal verification by automating theorem proving, generating invariants and specifications for complex systems, and improving solver efficiency and scalability. Humans: Frontier AI demonstrates limited capability in mitigating sophisticated human-targeted attacks (e.g., deepfakes), necessitating intensified research in this area.
5 Will Frontier AI Benefit Attackers or Defenders More?
@@
Building on our analysis of AI’s impact on cyber attacks and defense, this section predicts whether frontier AI will likely benefit attackers or defenders more in the short term. We base our analysis on: 1) equivalence class, where new defenses may equally benefit attackers, 2) the fundamental asymmetry between attackers and defenders, and 3) economic perspectives.
@@
TABLE IV: Equivalence classes: A list of defense and normal capabilities that can also help attacks.
@@
Proactive testing Pen. testing Enable more targeted attacks
@@
Attack detection ML-based threat detection Develop stronger evasion methods
@@
Triage forensic PoC & root cause Facilitate localization & exploitation
@@
Remediation Patch & testing generation Malware & weapon & exploit generation
@@
Normal utilities Multimodal generation Automated reconnaissance and delivery

### v3

@@ 4.3 Proactive and Reactive Defenses
Defense Steps 4&5: Remediation development and deployment. Research demonstrates frontier AI’s capability to automatically generate security vulnerability patches [133, 139, 126, 3, 276, 157, 177, 214, 263, 41]. To enhance patching effectiveness, researchers have explored various prompt strategies and built AI agents where LLMs can retrieve contextual information (e.g., caller functions of vulnerable code) [28, 139, 167, 265, 132]. The AIxCC competition demonstrates real-world applications, including patching an Off-by-One bug in SQLite3 [93]. However, no specific work exists on AI-facilitated remediation deployment.
@@ 4.4 Defenses with Provable Guarantees
Formal verification proves whether systems satisfy specific properties. This process is labor-intensive; seL4 required tens of proof-engineer years [102, 103]. Failed verifications require expert refinement, and software updates demand rewriting proofs, significantly increasing human effort. Recent research explores frontier AI to improve formal verification’s automation and scalability. Some works enhance interactive theorem proving [111, 264, 9, 258], e.g., GamePad [111] uses deep learning to synthesize Coq proofs. Others focus on automated program verification, where the typical solution is to model programs in domain-specific formal languages (DSLs), formulate security properties as specifications, and use constraint solvers for verification. Another line of research uses LLMs to generate program invariants, lifting a program to a DSL or supplementary text hints [178, 120, 26, 246, 148, 145]. For example, Loopy [120] leverages LLMs to generate candidate loop invariants and refined invariants via an SMT solver. Lemur [246] uses LLMs to generate high-level program invariants as sub-goals, which are then checked by solvers. LLMLift [26] combines LLMs with verified lifting tools to provide verified code transpilation from the source to DSLs, where LLMs generate lifted codes and proofs for functional equivalence. DafnyBench [148] evaluates SOTA LLMs in generating hints for the Dafny verifier. Other works improve solvers, including using traditional deep learning to enhance SAT and SMT solvers [21, 134] and leveraging LLMs for solver testing and fuzzing [219].
@@ 4.5 Defense For Humans
LLMs show promise in detecting social bots [74], and AI has been deployed for real-world fraud detection [131, 268, 60]. Research demonstrates multi-modal agents capable of detecting phishing websites [31]. AI-driven solutions are increasingly being applied to real-world scenarios. For example, one company deployed an AI-based honeypot to combat vishing, using a chatbot with a “grandma” voice to occupy fraudsters and waste their time [158]. AI-based deepfake detection continues as an active research area [187], with exploration of LLMs’ capabilities in identifying fake images [5]. Another significant direction focuses on AI-powered misinformation detection [12, 61, 142], often employing LLMs to identify LLM-generated content. Watermarking techniques represent another promising detection approach [277, 174, 141, 278, 87].
@@
Despite these active research efforts, Figure 2 shows current defensive techniques struggling to match sophisticated attacks. Research reveals significant limitations in AI systems’ defensive capabilities. For example, studies examining GPT-4’s ability to detect fake images [5] conclude the model lacks reliability and shows limitations with complex tasks. While many papers propose AI-powered defenses against deepfakes and social bots, most address traditional attack scenarios rather than novel or sophisticated attacks enabled by frontier AI. We need studies on adversarial dynamics between defense-based and attack-based AI systems—for example, how defensive AI can detect deepfakes or social bots generated by an AI attacker. This gap underscores the urgent need for research to develop AI-driven defenses against AI-driven human-targeted attacks.
5 Expert Survey
Expert surveys are a widely used approach for forecasting AI development timelines and risks [84, 217]. Accordingly, we conduct an expert survey to assess experts’ perceptions of frontier AI’s current impact on cybersecurity and their forecasts of its future impact.
(a) System attack steps
(b) Human-targeted attack activity
(c) Defense steps
Figure 3: Experts’ expected benefits of frontier AI across (a) system attack steps, (b) human-targeted attack activities, and (c) defense steps. In each plot, the $x$-axis represents the percentage of respondents selecting each benefit level.
5.1 Expert Views on AI’s Current Impact
Offensive Impact. We asked experts to rate the current effectiveness of AI in facilitating attacks on computer systems. Experts perceived frontier AI as moderately effective in accelerating such attacks (mean =3.2=3.2, std =1.0=1.0 on a 5-point scale, where higher values indicate greater effectiveness). Across seven attack steps, reconnaissance was viewed as the most influenced by frontier AI (mean =4.2=4.2, std =0.8=0.8), whereas access & remote control was rated least affected (mean =3.0=3.0, std =1.1=1.1) (Figure 3(a)).
The perceived effectiveness in accelerating human-targeted attacks was significantly higher than for system-targeted attacks (mean =3.7=3.7, std =1.0=1.0; $t=2.214$, $p=0.029$), which is consistent with our qualitative analysis in Section 4. Among the four human-targeted attack categories—social engineering, malicious content delivery, impersonation, and platform/service fraud—impersonation was rated highest (mean =4.5=4.5, std =0.7=0.7), while malicious content delivery was rated lowest (mean =3.7=3.7, std =1.2=1.2) (Figure 3(b)).
Defensive Impact. We also asked experts to evaluate how effective frontier AI is currently at helping to build defenses for computer systems. They rated its effectiveness between “slightly effective” and “moderately effective” (mean =2.7=2.7, std =0.9=0.9 on a 5-point scale). This rating is significantly lower than their perceived effectiveness for attacks (t=2.430t=-2.430, $p=0.017$). The lower ratings extend to defenses for humans (mean =2.4=2.4, std =0.9=0.9), which are significantly lower than for human-targeted attacks (t=6.295t=-6.295, $p=1.237\times 10^{-8}$). Thus, experts’ perceptions also reflect the relatively limited impact of frontier AI on defense.
Across the five defensive steps, consistent with our qualitative analysis in Section 4, remediation deployment received the lowest expected benefit from frontier AI (mean =3.1=3.1, std =1.1=1.1). By contrast, experts judged proactive testing and attack detection to be the most affected steps (proactive testing: mean =3.7=3.7, std =1.0=1.0; attack detection: mean =3.8=3.8, std =0.9=0.9), again aligning with our literature review analysis. Figure 3(c) presents the experts’ expected benefits for defenses.
5.2 Expert Views on Future AI Impacts
When Will AI Automate Attacks? We asked experts to estimate when AI could autonomously conduct eight system-targeted attack activities, using a six-point scale (“I believe AI can already do this”, “In the next two years”, “In 3–5 years”, “In the next 6–10 years”, “Longer than 10 years”, “Never”). On average, respondents placed these timelines just beyond the two-year mark (mean =2.3=2.3, std =1.1=1.1). Notably, many judged that AI can already autonomously craft phishing links and induce users to click malicious URLs (mean =1.4=1.4, std =0.6=0.6). By contrast, reverse-engineering complex software from stripped binaries was expected to take the longest (mean =2.7=2.7, std =1.3=1.3). Nevertheless, across all eight activities, the expected timelines fell within five years.
When asked about timelines for human-targeted attacks, experts projected shorter horizons than for system-targeted attacks (mean =2.0=2.0, std =1.2=1.2; t=1.788t=-1.788, $p=0.076$). Among the three human-targeted activities we evaluated, many experts judged that AI can already autonomously exploit deepfake technology and execute sophisticated business email compromise for fraud (deepfakes: mean =1.7=1.7, std =1.1=1.1; business email compromise: mean =1.7=1.7, std =0.8=0.8). By contrast, replicating biometric patterns (e.g., gait, typing patterns, behavioral biometrics) was assigned the longest timeline, falling between “in the next two years” and “in 3–5 years” on our scale (mean =2.6=2.6, std =1.4=1.4).
When Will AI Automate Defenses? Similarly, we estimated timelines for nine system defensive activities using the same six-point scale. Experts’ average estimates fell between “in the next two years” and “in 3–5 years” (mean =2.8=2.8, std =1.4=1.4). This estimate is significantly longer than the attack-side timeline ($t=3.892$, $p<0.001$), consistent with frontier AI’s stronger current impact on offense. Among the nine defensive activities, many experts judged that AI can already autonomously detect and block phishing messages containing malicious payloads with low false-positive and false-negative rates (mean =1.7=1.7, std =1.1=1.1). By contrast, fully automating root-cause analysis for all vulnerabilities—including N-day and zero-day—was expected to arrive later, trending toward the 6–10 year range (mean =3.4=3.4, std =1.4=1.4).
For the three human-focused defensive activities, expected timelines likewise fell between “in the next two years” and “in 3–5 years” (mean =2.5=2.5, std =1.3=1.3). Among the three activities, detecting and preventing sophisticated business email compromise was expected to arrive earliest (mean =2.2=2.2, std =1.4=1.4), whereas detecting and blocking unauthorized access attempts using synthetic or replicated biometric data was expected to take longest (mean =2.8=2.8, std =1.3=1.3). Consistent with the systems results, the expected timeline for human defenses was significantly longer than that for human-targeted attacks ($t=2.425$, $p=0.017$). Taken together, these findings indicate that experts expect frontier AI to enable autonomous attacks sooner than defenses.
Benefit Attackers or Defenders More? Cybersecurity’s dual nature naturally raises a central question: who will frontier AI benefit more, and how might this change over time? To probe views within the AI and security communities, we asked experts to assess how frontier AI will affect the overall balance between attackers and defenders across three timeframes—2, 5, and 10 years. They rated this balance on a 0-100 scale, where 100 indicates defenders exclusively benefit from AI, 50 indicates parity, and 0 indicates attackers exclusively benefit. Overall, experts anticipated that frontier AI would tilt the balance toward attackers in the near term, but that this advantage would narrow over time, approaching parity by the 10-year horizon (2 years: mean =34.2=34.2, std =25.3=25.3; 5 years: mean =43.6=43.6, std =21.2=21.2; 10 years: mean =49.8=49.8, std =26.0=26.0).
The most commonly mentioned reason that frontier AI will benefit attackers more is fundamental asymmetry; as one respondent put it, “In cybersecurity, there is a fundamental asymmetry: attackers need to find only one exploit, whereas defenders need everything to be secure. As such, I expect capability improvements to help attackers more first.” Moreover, defenders’ tolerance for failure is significantly lower than that of attackers, since failure can lead to substantial costs. As one respondent noted, “…defenders prioritize reliability and may take longer to trust and adopt AI tools in critical workflows.”
Economic rationality drives both attackers and defenders—attackers strike only when expected profits exceed costs, while defenders implement protections only when benefits outweigh costs. In line with this, strong economic incentives for attackers were mentioned as a reason: “Attackers can directly profit from intrusions, as evidenced by the rise of deepfake-related black markets.” Meanwhile, defenders often see less immediate financial return and face conflicts of interest: “Even if you have autonomous AI that can deploy defensive measures, they will likely incur some cost or run counter to some other system/organization stakeholder’s interests.”
However, many experts expected the gap to narrow over time. One reason that they cited is that defenders will possess greater resources, such as “vastly more legitimate data for training AI models” and “sustained enterprise investment in AI defensive capabilities”, than attackers. Priority access to advanced techniques (e.g., early-access programs for safety researchers) could further tilt the balance toward defense. Moreover, experts noted that defensive AI systems will learn from billions of real-world interactions with attackers, steadily increasing robustness. As defenses harden and systems become more secure, attackers must invest substantially more resources to discover vulnerabilities and conceal their identities, rendering many attacks impractical or cost-prohibitive.
Although most experts believed that frontier AI will benefit attackers more over the next two years as described above, about $29\%$ still expected the balance to tilt toward defenders (Figure 4 in Appendix). One respondent argued that “the current trajectory of frontier AI systems aligns more closely with defender use cases where any application is feasible.” Another one noted that “AI defenders can democratize the easier side of security, which is what most attacks target,” that is, by automating security hygiene (e.g., phishing filtering, rapid patching of known vulnerabilities, and misconfiguration remediation) and making default-secure configurations widely accessible. These divergent views underscore the uncertainty in predicting AI’s net impact on cybersecurity.
6 We Predict that AI will Benefit Attackers in the Short Term
@@
Similar to many experts, we argue that frontier AI will likely benefit attackers in the short term. In this section, we provide support based on two factors: 1) the equivalence class, where new defenses may equally benefit attackers, and 2) the fundamental asymmetry between attackers and defenders. We also discuss how this balance may evolve over time.
@@
Table 3: Equivalence classes: A list of defense and general capabilities that can also help attacks.
@@
Proactive
testing
Pen. testing Enable more targeted attacks
@@
Attack
detection
ML-based threat detection Develop stronger evasion methods
@@
Triage
forensic
PoC & root cause Facilitate localization & exploitation
@@
Remediation
Patch & testing generation Malware & weapon & exploit generation
@@
General use Multimodal generation Automated reconnaissance and delivery